A square metal plate with a length of 0.20 m and with a charge of 9.7 x 10 -7 C.


Find the electric field of the square metal plate.

1
Expert's answer
2022-02-17T04:36:20-0500

The electric field of a charged plate

"E=\frac{\sigma}{2\epsilon_0}=\frac{Q}{2\epsilon_0 A}"

"E=\frac{9.7* 10^{ -7}}{2*8.85*10^{-12}*(0.20)^2}=1.37*10^6\:\rm V/m"